OpenStudy (saucvi):

Question about system of equations: for equivalent system of equations if both equations start with a negative, what do you multiply? Do the numbers stay negative?

OpenStudy (saucvi):

Example: -3x + 7y =-16 -9x + 5y = 16 how would you get an equivalent system of equations for that?

OpenStudy (mertsj):

Multiply the top equation by -3

OpenStudy (saucvi):

9x + -27y = 48 ?

OpenStudy (mertsj):

-3(7)=-21

OpenStudy (mertsj):

Better learn you multiplication tables, sonny.

OpenStudy (saucvi):

I'm not understanding this at all.

OpenStudy (mertsj):

What don't you understand about multiplying 3and 7 to get 21 and NOT 27?

OpenStudy (saucvi):

I don't understand why you multiplied 7 by -3 in the first place. I thought you multiply each variable to get an equivalent.

OpenStudy (mertsj):

To solve a system of equations you want to make the coefficients of either x or y to be opposites.

OpenStudy (mertsj):

If you multiply the first equation by -3 you will get 9x.

OpenStudy (mertsj):

In the second equation you have -9x

OpenStudy (mertsj):

9 and -9 are opposites.

OpenStudy (saucvi):

Oh okay. I think I understand most of it now. Thanks. I appreciate it.